OpenStudy (anonymous):

Will give medal~ Can someone check my answer? It's a volume question... The volume of a cylinder is 24 ft^3. The volume of a cube is 54 ft^3. What is the difference between the volume of a cone that fits exactly inside the cylinder and the volume of a pyramid that fits exactly inside the cube? Provide an explanation for your answer to receive full credit.

OpenStudy (anonymous):

So the volume for the cone would be 1/3 the volume of the cylinder (from the formula 1/3πr^2). The same goes for the pyramid inside the cube. Volume of cone: = 1/3 * 24 = 8ft^3 Volume of pyramid: = 1/3 * 54 = 18ft^3 So the difference between the cone and the pyramid is 10ft^3

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

Thank ya kindly